OpenStudy (anonymous):

PLEASE HELP!! SHOW ALL WORK Write a polynomial function with rational coefficients so that P(x)=0 has the given roots. 1. 3i2 2. 1,2,-i 3.2+squareroot ,3 4. 2,4+i THANKS A LOT!!!!

OpenStudy (anonymous):

what is the first one?

OpenStudy (anonymous):

just \(3i\) ?

OpenStudy (anonymous):

ok that is the easiest of the bunch if \(3i\) is a zero, then so is \(-3i\) and in factored form the polynomial will be \[P(x)=(x+3i)(x-3i)\]

OpenStudy (anonymous):

when you multiply that out you get \[P(x)=x^2+9\] which you might have guessed at to begin with since if you go to solve \[x^2+9=0\] you get \[x^2+9=0\\ x^2=-9\\ x=\pm 3i\]

OpenStudy (anonymous):

\(1,2,-i\) is similar if \(-i\) is a zero then so is \(i\) so you have in factored form \[P(x)=(x-1)(x-2)(x-i)(x+i)\] and you multiply that mess out

OpenStudy (anonymous):

not as bad as it looks, since \((x-i)(x+1)=x^2+1\) so your real job is to try and multiply \[(x-1)(x-2)(x^2+1)\] with out making an algebra mistake i would cheat

OpenStudy (anonymous):

how do you do it? @satellite73

OpenStudy (anonymous):

how do i multiply it out?

OpenStudy (anonymous):

can you tell me the answers to all the questions too? plzzz @satellite73

OpenStudy (anonymous):

not sure what the question is is it how do i multiply out \((x-1)(x-2)(x^2+1)\) ? besides cheating, i would start with \[(x-1)(x-2)=x^2-3x+2\] and then multiply \[(x^2-3x+2)(x^2+1)\] using six multiplications

OpenStudy (anonymous):

can you do that or not?

OpenStudy (anonymous):

what about the others?

OpenStudy (anonymous):

no im confused and i need to bring the grade up tomoorow is the last day to turn it in

OpenStudy (anonymous):

is the third one \(2+\sqrt3\) ?

OpenStudy (anonymous):

i really need a good grade i have an F :(

OpenStudy (anonymous):

you don't want an F that is for sure

OpenStudy (anonymous):

[2+sqrt{5} 3

OpenStudy (anonymous):

yah i dont thats why i need to get this done so atleast it goes to a D or a C

OpenStudy (anonymous):

ok lets take is slow and finish the first one if you want to multiply \((x-1)(x-2)(x^2+1)\) and make sure you don't get a mistake (which is also making sure you don't actually learn how to do it, but whatever) use this http://www.wolframalpha.com/input/?i=%28x-1%29%28x-2%29%28x^2%2B1%29

OpenStudy (anonymous):

the second one i take it is \(2+\sqrt5,3\) is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that site is so confusing

OpenStudy (anonymous):

then if \(2+\sqrt5\) is a zero, so is \(2-\sqrt5\) oh no the answer is right there right where it says "alternate form" do you see it?

OpenStudy (anonymous):

you get two graphs, and right underneath it says "alternate form" which means they multiplied it out for you and you see \[x^4-3 x^3+3 x^2-3 x+2\]which is your answer

OpenStudy (anonymous):

ready for the one with zeros \(2+\sqrt5,2-\sqrt5,3\) ?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

number 3 it says the zeros are \(2+\sqrt5\) and \(3\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so you know one factor is \((x-3)\) since \(3\) is a zero your hard job is to come up with the quadratic whose zero is \(2+\sqrt5\) but there is an easy way to do this do you know it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

set \[x=2+\sqrt5\] then subtract \(2\) from both sides and get \[x-2=\sqrt5\] then square both sides you get \[(x-2)^2=5\]or \[x^2-4x+4=5\] and finally subtract \(5\) from both sides to get \[x^2-4x-1=0\] so your quadratic is \[x^2-4x-1\] and you final job is to multiply \[(x^2-4x-1)(x-3)\]

OpenStudy (anonymous):

how to multiply?

OpenStudy (anonymous):

take \(x^2-4x-1\) and multiply each term by \(x\) then multiply each term by \(-3\) then combine like terms

OpenStudy (anonymous):

you know how to do that?

OpenStudy (anonymous):

so than what is the answers?

OpenStudy (anonymous):

if you just want the answer, do this http://www.wolframalpha.com/input/?i=%28x^2-4x-1%29%28x-3%29 and read where it says "alternate form"

OpenStudy (anonymous):

there is no alternate form for number 4

OpenStudy (anonymous):

lol we didn't do number 4 yet that is \(2,4+i\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok for this one set \[x=4+i\] subtract 4 get \[x-4=i\] square and get \[x^2-8x+16=-1\] add \(1\) to get \[x^2-8x+17\] then multiply \[(x^2-8x+17)(x-2)\]

OpenStudy (anonymous):

u know i just want to get this done and its getting late and i have to sleep and i really appreciate u helping me with this! ur the only one THANKS

OpenStudy (anonymous):

alternate form says \(x^3-10 x^2+33 x-34\) for that one http://www.wolframalpha.com/input/?i=%28x^2-8x%2B17%29%28x-2%29 you got them all now?

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

thanks alot ur great and im saying it in a good way not a bad way

OpenStudy (anonymous):

thanks my pleasure hope you have all the right answers and don't get an F

OpenStudy (anonymous):

there is something more if u can help?

OpenStudy (anonymous):

sure go ahead and ask

OpenStudy (anonymous):

Find the real and/or imaginary solution of each polynomial equation 1. \[4x^3+4=0\] 2. x^4-x^2-72=0 3.8x^3=-1 4. x^3=64 5.8x^3+27=0 6.2x^4+16x^2=40 7.2x^4-x^2=3 8.x^4-7x^2-8=0

OpenStudy (anonymous):

first one divide by 4 and start with \[x^3+1=0\] then factor as \[(x+1)(x^2-2x+1)=0\]so \(x=-1\) or \(x^2-2x+1=0\) which does not factor, so you need to use the quadratic formula

OpenStudy (anonymous):

unless it was \(4x^3+4x=0\) but i assume what you wrote was correct

OpenStudy (anonymous):

i made a typo there, it should be \((x+1)(x^2-x+1)=0\)

OpenStudy (anonymous):

there is no x after the second 4

OpenStudy (anonymous):

the zeros of \(x^2-x+1\) are \(\frac{1\pm\sqrt{3} i}{2}\)

OpenStudy (anonymous):

so you have 3 zeros: \(-1,\frac{1+\sqrt{3}i}{2},\frac{1-\sqrt{3}i}{2}\)

OpenStudy (anonymous):

second one factors \[x^4-x^2-72=(x^2-9)(x^2+8)\]

OpenStudy (anonymous):

set \(x^2-9=0\) get \((x-3)(x+3)=0\) so \(x=3\) or \(x=-3\)

OpenStudy (anonymous):

set \(x^2+8=0\) get \(x^2=-8\) so \(x=\pm\sqrt{8}i\)

OpenStudy (anonymous):

or if you prefer \(x=\pm2\sqrt{2}i\)

OpenStudy (anonymous):

\(8x^3=-1\) gives \[x^3=-\frac{1}{8}\]so \(x=-\frac{1}{2}\)

OpenStudy (anonymous):

oh you need the complex zeros too?

OpenStudy (anonymous):

idk it doesnt say